# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

### v5

@@ 2.1 Tempered and discrete series representations
A smooth representation $\pi$ of $G$ is tempered if $\pi$ is unitary, of finite length, and all matrix coefficients of $\pi$ belong to $L^{2+\epsilon}(G/Z(G))$ for all $\epsilon>0$. We write $\mathcal{M}_{t}(G)$ for the category of tempered representations of $G$.
In particular, the central character of such a representation takes values in the circle group $\mathbb{T}\subset\mathbb{C}^{\times}$.
@@
A smooth finite-length representation $\omega$ of $G$ belongs to the discrete series if all matrix coefficients of $\omega$ are square-integrable modulo $Z(G)$. We write $\mathcal{E}_{2}(G)$ for the space of unitary discrete series, and $\mathcal{E}_{2}(G)^{I}$ for the space of unitary discrete series with nontrivial Iwahori-fixed vectors.
@@ 2.1.1 Formal degrees of discrete series representations
This result is proven by first proving that the Hiraga-Ichino-Ikeda conjecture [HII] holds for unipotent discrete series representations. Note that [Solleveld], [fomFormalDegUnip] and [HII] all use the normalization of the Haar measure on $G$ defined in [HII]. This normalization gives in our setting $\mu_{\mathrm{HII}}(K)=q^{\dim\mathbf{G}}\#\mathbf{G}(\mathbb{F}_{q})$. Hence, noting that #$\#\mathbf{G}(\mathbb{F}_{q})=P_{\mathbf{G}/\mathbf{B}}(q)\cdot\#\mathbf{B}(\mathbb{F}_{q})$ and that, as $\mathbb{F}_{q}$ is perfect, #$\#B(\mathbb{F}_{q})$ is a polynomial in $q$, we have
@@ 2.2.1 Unramified characters
Let $X(\mathbf{G})=\mathrm{Hom}(\mathbf{G},{\mathbb{G}_{\mathrm{m}}})$ denote the rational characters of $\mathbf{G}$ defined over $F$. Let $\mathfrak{a}_{G}:=(X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{R})^{*}=(X^{*}(\mathbf{G})\otimes_{\mathbb{Z}}\mathbb{R})^{*}$ be the real Lie algebra of the maximal split central torus $\mathbf{A}=\mathbf{A}_{G}$ of $\mathbf{G}$ and let ${\mathfrak{a}_{G}}_{\mathbb{C}}$ be its complexification. We have a map
$X(\mathbf{G})\to\mathrm{Hom}(G/G^{1},\mathbb{C}^{\times})=:\mathcal{X}(G)$
given by $\chi\mapsto|\chi|_{F}$, where $|\chi|_{F}(g)=|\chi(g)|_{F}$ and $G^{1}=\bigcap_{\chi\in X(\mathbf{G})}\ker|\chi|_{F}$. This gives the unramified characters $\mathcal{X}(G)$ a complex manifold structure under which $\mathcal{X}(G)\simeq(\mathbb{C}^{\times})^{\dim_{\mathbb{R}}\mathfrak{a}_{G}}$. For indeed, we have
${{\mathfrak{a}_{G}}_{\mathbb{C}}^{*}}$${\mathcal{X}(G)}$11{1}1
@@ 2.2.2 Unitary unramified characters
Denote $\mathrm{Im}\,\mathcal{X}(G)$ the set of unitary unramified characters taking values in the unit circle $\mathbb{T}\subset\mathbb{C}^{\times}$. This notation is justified, as if $|\chi|_{F}=|\chi^{\prime}|_{F}$, then $\mathrm{Re}(\chi)=\mathrm{Re}(\chi^{\prime})\in\mathfrak{a}_{G}$. Hence we can define $\mathrm{Im}\,\mathcal{X}(G)$ to be the unramified characters coming from pure imaginary elements of ${\mathfrak{a}_{G}}_{\mathbb{C}}$.
@@ 2.2.3 Action by twisting and the canonical measure
The group $\mathcal{X}(G)$ acts on admissible representations of $G$ by twisting: $\omega\mapsto\omega\otimes\nu$ for $\nu\in\mathcal{X}(G)$. This restricts to an action the unitary unramified characters on $\mathcal{E}_{2}(G)$.
Pulling this action back to $i\mathfrak{a}_{G}$, and given a representation $\omega$, let $L^{*}$ be its stabilizer in $i\mathfrak{a}_{G}$, so that the orbit $\mathfrak{o}$ of $\omega$ is identified as $i{\mathfrak{a}_{G}}/L^{*}$. This gives $\mathfrak{o}=\mathrm{Im}\,\mathcal{X}(G)\cdot\omega$ the structure of a real submanifold of the larger orbit $\mathfrak{o}_{\mathbb{C}}={\mathfrak{a}_{G}}_{\mathbb{C}}/L^{*}=\mathcal{X}(G)\cdot\omega$.
@@
If $G=\mathrm{SL}_{2}(F)$ and $M=A$ is the diagonal torus, we have $\mathfrak{a}_{G}^{*}\simeq\mathbb{R}$ and $R=\mathbb{Z}$ so that a fundamental domain for $\mathfrak{a}_{G}^{*}/\frac{2\pi}{\log q}R$ is $\left[-\frac{\pi}{\log q},\frac{\pi}{\log q}\right)$ and the canonical measure $d\nu=\frac{\log q}{2\pi}\mathrm{d}x$, where $\mathrm{d}x$ is the Lebesgue measure. To obtain quasicharacters of $G$, we associate to $\nu\in{\mathfrak{a}_{G}}_{\mathbb{C}}$ the quasicharacter $\chi_{\nu}(g)=q^{i\left\langle\nu,H_{G}(g)\right\rangle}$.
@@
In general, $\mathcal{E}_{2}(M_{P})$ is a disjoint union of compact tori, and the Plancherel density descends to the quotients of these compact tori by certain finite groups, namely, the Weyl groups of $(P,A_{P})$. The set $\mathcal{E}_{2}(M_{P})^{I}$ is finite up to twist by unramified characters, by a result of Harish-Chandra [Waldspurger].
@@ 2.3 The Harish-Chandra Schwartz algebra
Define the subring $\mathcal{E}(G)$ of the endomorphism ring of the forgetful functor $\mathcal{M}(G)\to\mathbf{Vect}_{\mathbb{C}}$ from the category of all smooth representations, by replacing, in (1) above, $\mathcal{E}_{2}(M)$ with $\mathcal{M}(M)$, unitary characters with all unramified characters, and “smooth” with ”algebraic.” The matrix Paley-Wiener theorem of Bernstein [Bernstein] says that $f\mapsto\pi(f)$ is an isomorphism from the full Hecke algebra of $G$ onto $\mathcal{E}(G)$. Denote $\mathcal{E}^{I}$ and $\mathcal{E}_{t}$ the subrings of $I$-invariant endomorphisms.
@@
We say a non-strictly positive character $\chi$ is strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x)|_{\infty}>1$ for $|x|_{F}\geq 1$.
Of course, it suffices to test this for $x=\varpi^{-1}$. The condition for $\chi^{-1}$ to be non-strictly positive is then that $|\chi(\alpha^{\vee}(\varpi))|_{\infty}\geq 1$.
@@
For $\mathbf{G}=\mathrm{SL}_{2}$, in the conventions fixed in Example 2, an unramified character $\chi$ of $A$ is non-strictly positive if $\chi^{-1}$ corresponds to $z$ such that $|z|\geq 1$.
If $\mathbf{G}=\mathrm{GL}_{n}$ and $\nu$ corresponds to the vector $(z_{1},\dots,z_{n})\in\mathbb{C}^{n}$, then the condition that $\nu^{-1}$ is non-strictly positive translate to $|z_{n}|\geq|z_{n-1}|\geq\cdots\geq|z_{1}|$. Such conditions divide $\mathbb{C}^{n}$ into chambers, on which the Weyl group $\mathfrak{S}_{n}$ clearly acts simply-transitively. Interior points correspond to $\nu$ such that $\nu^{-1}$ is strictly positive.
@@
    For all $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$, the endomorphism $\eta_{\pi}=\eta_{\nu,\omega}$ is a rational function of $\nu$, regular on the set of $\nu$ such that $\nu^{-1}$ is non-strictly positive.
@@
    Let $\pi$ be a tempered representation of $G$. Then the action of $H$ on $\pi^{I}$ extends uniquely to $J$.
@@
    Let $P=MN$ be a parabolic subgroup of $G$ with Levi subgroup $M$ and let $\omega$ an irreducible tempered representation of $M$. Let $\nu$ be a non-strictly positive character of $M$ and let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu^{-1})$. Then the action of $H(G,I)$ on $\pi^{I}$ extends uniquely to an action of $J$.
@@
    The map $t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}$ defines a homomorphism
@@
Composing this isomorphism with the inverse Fourier transform, Braverman and Kazhdan define an algebra map
@@
$t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}\mapsto f_{w}=\sum_{x\in\tilde{W}}A_{x,w}T_{x}\in\mathcal{C}^{I\times I},$
@@
Let $\pi$ be a tempered representation of $G$. By [BK], the $H$-action on it extends to a $J$-action, and by Corollary 2.6 of loc. cit., we have $\eta_{\pi}(\phi_{q}(f^{\dagger}))=\pi(f)$ in $\mathcal{E}_{t}^{I}$ for any $f\in H$. Therefore $\tilde{\phi}\circ\phi_{q}(f^{\dagger})=f$ by the Plancherel theorem. ∎
Remark 6.
The presence of the Goldman involution is necessary for tempered representations of $J$ to restrict to tempered representations of $H$. For example, in [BK], Braverman and Kazhdan showed for $G=\mathrm{SL}_{2}(F)$, that $\mathcal{E}^{I}_{J}(G)=\operatorname{End}(\mathrm{St}^{I})\oplus\operatorname{End}_{\Phi,X_{*}}(\mathcal{S}^{I})$, matching the decomposition $J=\mathbb{C}t_{1}\oplus J_{0}$, where $\Phi$ is the Fourier transform on the Schwartz space of the basic affine space $\mathcal{S}$ [BKBasicAffine], and $\mathrm{St}$ is the Steinberg representation. In this case $J$ has three one-dimensional representations, whereas $H$ has one one-dimensional discrete series representation $\mathrm{St}$, two tempered one-dimensional representations not belonging to the discrete series, and one non-tempered one-dimensional representation, the trivial representation. It is easy to compute using the formulas of [D] that $\phi_{q}(T_{s})t_{1}=qt_{1}$, whereas $\phi_{q}(T_{s}^{\dagger})t_{1}=-t_{1}$. Thus for the Steinberg representation to restrict to the Steinberg representation, the twist is required.

### v6

@@ 2.1 Tempered and discrete series representations
We take our definitions of discrete series and tempered representations from [Waldspurger, III.1] and [Waldspurger, III.2] respectively. By parabolic induction we always understand normalized induction.
@@
A smooth admissible representation $\omega$ of $G$ belongs to the discrete series if $\omega$ admits a unitary central character and all matrix coefficients of $\omega$ are square-integrable modulo $Z(G)$. We write $\mathcal{E}_{2}(G)$ for the space of irreducible discrete series, and $\mathcal{E}_{2}(G)^{I}$ for the space of irreducible discrete series with nontrivial Iwahori-fixed vectors.
Let $v$ be the $K$-fixed vector in the self-contragredient representation $\operatorname{Ind}_{B}^{G}(\mathrm{triv})^{K}$ such that $v(1)=1$. Define $\Xi(g)=\left\langle\pi(g)v,v\right\rangle$ to be the corresponding matrix coefficient.
@@
A smooth function $f$ on $G$ is tempered if there is $C>0$ and $r\in\mathbb{R}$ such that
$|f(g)|\leq C\Xi(g)\left(1+\log\left\|g\right\|\right)^{r},$
where $\left\|g\right\|\geq 1$ is defined as in [Waldspurger, p.242].
Definition 9.
A smooth admissible representation $\pi$ of $G$ is tempered if all its matrix coefficients are tempered functions in the sense above.
We write $\mathcal{M}_{t}(G)$ for the category of tempered representations of $G$. If a tempered representation admits a central character, the central character takes values in the circle group $\mathbb{T}\subset\mathbb{C}^{\times}$.
@@ 2.1.1 Formal degrees of discrete series representations
This result is proven by first proving that the Hiraga-Ichino-Ikeda conjecture [HII] holds for unipotent discrete series representations. Note that [Solleveld], [fomFormalDegUnip] and [HII] all use the normalization of the Haar measure on $G$ defined in [HII]. This normalization gives in our setting $\mu_{\mathrm{HII}}(K)=q^{\dim\mathbf{G}}\#\mathbf{G}(\mathbb{F}_{q})$. Hence, noting that $\mathbf{G}(\mathbb{F}_{q})=P_{\mathbf{G}/\mathbf{B}}(q)\cdot\#\mathbf{B}(\mathbb{F}_{q})$ and that, as $\mathbb{F}_{q}$ is perfect, $B(\mathbb{F}_{q})$ is a polynomial in $q$, we have
@@ 2.2.1 Unramified characters
Let $X^{*}(\mathbf{G})=\mathrm{Hom}(\mathbf{G},{\mathbb{G}_{\mathrm{m}}})$ denote the rational characters of $\mathbf{G}$ defined over $F$. Let $\mathfrak{a}_{G}:=(X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{R})^{*}=(X^{*}(\mathbf{G})\otimes_{\mathbb{Z}}\mathbb{R})^{*}$ be the real Lie algebra of the maximal split central torus $\mathbf{A}=\mathbf{A}_{G}$ of $\mathbf{G}$ and let $\mathfrak{a}_{G}$$_{\mathbb{C}}$ be its complexification. We have a map
$X^{*}(\mathbf{G})\to\mathrm{Hom}(G/G^{1},\mathbb{C}^{\times})=:\mathcal{X}(G)$
given by $\chi\mapsto|\chi|_{F}$, where $|\chi|_{F}(g)=|\chi(g)|_{F}$ and $G^{1}=\bigcap_{\chi\in X^{*}(\mathbf{G})}\ker|\chi|_{F}$. This gives the unramified characters $\mathcal{X}(G)$ a complex manifold structure under which $\mathcal{X}(G)\simeq(\mathbb{C}^{\times})^{\dim_{\mathbb{R}}\mathfrak{a}_{G}}$. For indeed, we have
$\mathfrak{a}_{G}$$_{\mathbb{C}}^{*}$$\mathcal{X}(G)$}1{1}
@@ 2.2.2 Unitary unramified characters
Denote $\mathrm{Im}\,\mathcal{X}(G)$ the set of unitary unramified characters taking values in the unit circle $\mathbb{T}\subset\mathbb{C}^{\times}$. This notation is justified, as if $|\chi|_{F}=|\chi^{\prime}|_{F}$, then $\mathrm{Re}(\chi)=\mathrm{Re}(\chi^{\prime})\in\mathfrak{a}_{G}^{*}$. Hence we can define $\mathrm{Im}\,\mathcal{X}(G)$ to be the unramified characters coming from pure imaginary elements of $\mathfrak{a}_{G}$$_{\mathbb{C}}^{*}$.
@@ 2.2.3 Action by twisting and the canonical measure
The group $\mathcal{X}(G)$ acts on admissible representations of $G$ by twisting: $\omega\mapsto\omega\otimes\nu$ for $\nu\in\mathcal{X}(G)$. This restricts to an action of $\mathrm{Im}\,\mathcal{X}(G)$ on $\mathcal{E}_{2}(G)$.
Pulling this action back to $i\mathfrak{a}_{G}^{*}$, and given a representation $\omega$, let $L^{*}$ be its stabilizer in $i\mathfrak{a}_{G}^{*}$, so that the orbit $\mathfrak{o}$ of $\omega$ is identified as $i{\mathfrak{a}_{G}}^{*}/L^{*}$. This gives $\mathfrak{o}=\mathrm{Im}\,\mathcal{X}(G)\cdot\omega$ the structure of a real submanifold of the larger orbit $\mathfrak{o}_{\mathbb{C}}={\mathfrak{a}_{G}}^{*}_{\mathbb{C}}/L^{*}=\mathcal{X}(G)\cdot\omega$.
@@
If $G=\mathrm{SL}_{2}(F)$ and $M=A$ is the diagonal torus, we have $\mathfrak{a}_{G}^{*}\simeq\mathbb{R}$ and $R=\mathbb{Z}$ so that a fundamental domain for $\mathfrak{a}_{G}^{*}/\frac{2\pi}{\log q}R$ is $\left[-\frac{\pi}{\log q},\frac{\pi}{\log q}\right)$ and the canonical measure $d\nu=\frac{\log q}{2\pi}\mathrm{d}x$, where $\mathrm{d}x$ is the Lebesgue measure. To obtain quasicharacters of $G$, we associate to $\nu\in{\mathfrak{a}_{G}}^{*}_{\mathbb{C}}$ the quasicharacter $\chi_{\nu}(g)=q^{\left\langle\nu,H_{G}(g)\right\rangle}=|\nu(g)|_{F}$, where the second equality defines $H_{G}\colon G\to\mathfrak{a}_{G}$.
@@
In general, $\mathcal{E}_{2}(M_{P})$ is a disjoint union of compact tori, and the Plancherel density descends to the quotients of these compact tori by certain finite groups, namely, the Weyl groups of $(P,A_{P})$. The set $\mathcal{E}_{2}(M_{P})^{I}$ is finite up to twists by unramified characters, by a result of Harish-Chandra [Waldspurger].
@@ 2.3 The Harish-Chandra Schwartz algebra
Define the subring $\mathcal{E}(G)$ of the endomorphism ring of the forgetful functor $\mathcal{M}(G)\to\mathbf{Vect}_{\mathbb{C}}$ from the category of all smooth representations, by replacing, in (1) above, unitary characters with all unramified characters, and “smooth” with ”algebraic”, and adding
  1. 3.
    The endomorphisms $\eta_{\pi}$ are compatible with supercuspidal support, i.e. if $(M_{1},\sigma_{1})$ is the supercuspidal support of $\pi$, so that
    $\pi=i_{P}^{G}(\omega\otimes\nu)\mathrel{\hookrightarrow}\pi_{1}=i_{P_{1}}^{G}(\sigma_{1}\otimes\nu_{1}),$
    then $\eta_{\pi_{1}}$ preserves $\pi$ and $\eta_{\pi_{1}}|_{\pi}=\eta_{\pi}$.
The matrix Paley-Wiener theorem of Bernstein [Bernstein] says that $f\mapsto\pi(f)$ is an isomorphism from the full Hecke algebra of $G$ onto $\mathcal{E}(G)$. Denote $\mathcal{E}^{I}$ and $\mathcal{E}_{t}$ the subrings of $I$-invariant endomorphisms.
Remark 6.
Property 3 means that unlike Schwartz functions, the Fourier transform a compactly supported function may be freely specified only on inductions of supercuspidal representations. In particular, an algebraic family of endomorphisms defined for all unramified characters may fail to come from a compactly-supported function. An example of a non-compactly supported function with regular Fourier transform is the element of $J$ given in Corollary 4, because $\pi(t_{1})=0$ except if $\pi$ is the Steinberg representation of $G$.
@@
We say a non-strictly positive character $\chi$ is strictly positive if for all root subgroups $U_{\alpha}\subset N_{P}$, we have $|\chi(\alpha^{\vee}(x))|_{\infty}>1$ for $|x|_{F}>1$.
Of course, it suffices to test this for $x=\varpi^{-1}$.
@@
For $\mathbf{G}=\mathrm{SL}_{2}$, in the conventions fixed in Example 2, an unramified character $\chi$ of $A$ is non-strictly positive if it corresponds to $z$ such that $|z|\geq 1$.
If $\mathbf{G}=\mathrm{GL}_{n}$ and $\nu$ corresponds to the vector $(z_{1},\dots,z_{n})\in\mathbb{C}^{n}$, then the condition that $\nu$ is non-strictly positive translates to $|z_{1}|\geq|z_{2}|\geq\cdots\geq|z_{n}|$. Such conditions divide $\mathbb{C}^{n}$ into chambers, on which the Weyl group $\mathfrak{S}_{n}$ clearly acts simply-transitively. Interior points correspond to strictly positive $\nu$.
@@
    For all $\pi=\operatorname{Ind}_{P}^{G}(\nu\otimes\omega)$, the endomorphism $\eta_{\pi}=\eta_{\nu,\omega}$ is a rational function of $\nu$, regular on the set of non-strictly positive $\nu$.
@@
    Let $\pi$ be a tempered representation of $G$. Then the action of $H$ on $\pi^{I}$ extends uniquely to $J$ via $\phi\circ{}^{\dagger}(-)$.
@@
    Let $P=MN$ be a parabolic subgroup of $G$ with Levi subgroup $M$ and let $\omega$ an irreducible tempered representation of $M$. Let $\nu$ be a non-strictly positive character of $M$ and let $\pi=\operatorname{Ind}_{P}^{G}(\omega\otimes\nu)$. Then the action of $H(G,I)$ on $\pi^{I}$ extends uniquely to an action of $J$.
@@
    The action of $J$ on the representations $\pi^{I}$ in (2) extends rationally in $\nu$ to define a homomorphism
@@
    denoted
    $t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}.$
The proof of Theorem 7 (1) and (2) in [BK] uses Theorem 4 and [XiJAMS], to show that, in the notation of Section 1.5.1,
$E(u,s,\rho)|_{H}\simeq{}^{*}K(u,s,\rho,q)={}^{*}i_{P}^{G}(\sigma\otimes\nu)$
where the restriction is via $\phi$, and hence that
$E(u,s,\rho)_{H,\phi\circ{}^{*}(-)}\simeq K(u,s,\rho,q)=i_{P}^{G}(\sigma\otimes\nu),$
where the restriction is now via $\phi\circ{}^{*}(-)$.
By Lemma 1, the map $\phi$ is $\bar{\kappa}$-linear, where $\bar{\kappa}$ is as in Definition 6. Of course, $\phi$ also intertwines conjugation by $T_{w_{0}}$ and $\phi(T_{w_{0}})$.
As noted in Section 1.5.1, if $\pi=K(u,s,\rho)$ is simple tempered (or more generally, is any semisimple module), then by [PrasadDuke, Prop. 6.3],
$E(u,s,\rho)|_{\phi\circ(-)^{\dagger}}={}^{\dagger}{}^{*}K(u,s,\rho)\simeq\widetilde{K(u,s,\rho)}$ (4)
is simple tempered. Moreover, by Lemma 2 (e) and (4), we have, if $\Psi_{w_{0}}(h)=T_{w_{0}}hT_{w_{0}}^{-1}$, that
${}^{\bar{\kappa}\circ\phi(\Psi_{w_{0}}^{-1})}E(u,s,\rho)|_{\phi\circ{}^{\dagger}(-)}$ $={}^{\phi(\Psi_{w_{0}}^{-1})}E(u,s,\rho)|_{\bar{\kappa}\circ\phi\circ{}^{\dagger}(-)}$
$=E(u,s,\rho)|_{\phi(\Psi_{w_{0}}^{-1})\circ\bar{\kappa}\circ\phi\circ{}^{\dagger}(-)}$
$=E(u,s,\rho)|_{\phi\circ\Psi_{w_{0}}^{-1}\circ\bar{\kappa}\circ{}^{\dagger}(-)}$
$=E(u,s,\rho)|_{\phi\circ{}^{*}(-)}$
$=K(u,s,\rho)$
for any standard $H$-module $K(u,s,\rho)$. Therefore any standard $H$-module extends to a simple $J$-module via $\phi\circ{}^{\dagger}(-)$.
Composing the morphism $\eta$ with the inverse Fourier transform, Braverman and Kazhdan define an algebra map
@@
$t_{w}\mapsto(\eta_{\pi}(w))_{\pi\in\mathcal{M}_{t}(G)}\mapsto f_{w}=\sum_{x\in W_{\mathrm{aff}}}A_{x,w}T_{x}\in\mathcal{C}^{I\times I},$
@@
Let $\pi$ be a tempered representation of $G$. By Theorem 7 and (4), the $H$-action on it extends to a $J$-action via $\phi\circ^{\dagger}(-)$, such that
$\eta_{\pi}(\phi_{q}({}^{\dagger}f))=\pi(f)$
in $\mathcal{E}_{t}^{I}$ for any $f\in H$. Therefore $\tilde{\phi}\circ\phi_{q}({}^{\dagger}f)=f$ by the Plancherel theorem. ∎